# Describing Hadronization via Histories and Observables for Monte-Carlo Event Reweighting

### v1

@@ 2.1 Lund string fragmentation model for the $q\bar{q}$ case
where the term $bA$ corresponds to the imaginary part of the action of a massless string. The area $A$ is space-time area scaled by $\kappa^{2}$. Generating the transition from a string state to a hadronic state which fulfills eq. 1, is now a question of selecting a specific implementation. In the Lund model, the increasing separation between the quark and the antiquark, makes it energetically favorable to create $q_{j}\bar{q}_{j}$ pairs out of the vacuum. The string therefore breaks into spacelike separated fragments, i.e., hadrons. Due to the spacelike separation, there is no preferred time-ordering, and the hadronization process could therefore be described either by an inside-out cascade, starting from the center and fragmenting outwards, or an outside-in cascade, starting from the string ends. The Lund model makes the latter selection. Starting from one string end, a string break now produces a hadron containing the string end quark (or anti-quark), and the anti-quark (or quark) from the string break. In this way, the $j$’th string break will produce the hadron containing $q_{j-1}\bar{q}_{j}$ or vice versa. We call this sequence of multiple hadron emissions from a single string fragment a fragmentation chain.
@@
Iterative fragmentations continue until the energy of the string system crosses a chosen low-energy threshold. The remaining string piece is then combined into a final pair of hadrons. Within the Pythia event generator code, this final combination is performed by an algorithm called finalTwo. The finalTwo method effectively works as a filter by checking whether the final hadrons can be produced on-shell while ensuring that the overall fragmentations follow the left-right symmetric Lund fragmentation function. If this is not the case, the generated fragmentation chain is rejected, and the simulation of hadronization starts anew, taking again the original string as the starting point. This process can be repeated several times, until the finalTwo step is successful.
@@
String breaks are described by seven dimensional vectors containing: the light-cone momentum fraction of the hadron, $z$; the two-dimensional momentum kick $\Delta\rm=(\Delta p_{x},\Delta p_{y})$ of the emitted hadron; the hadron mass $m_{ij}$; a boolean that encodes whether the string break occurred at the positive or the negative end of the string, stored within Pythia as fromPos; and the transverse momentum of the string before the breaking, =$\rm{}^{\phantom{,}\text{string}}=(p_{x}^{\text{string}},p_{y}^{\text{string}})$ [^6] In general, the energy of the string in its center-of-mass frame, $E_{\text{string}}$, would also be included in this list. We are able to use the shorter list in eq. 4, since the fragmentations in the Lund string model do not depend explicitly on $E_{\text{string}}$. Additional information must also be included in the case where gluons are attached to the string, and the initial state changes between events.
@@
while a vector of rejected fragmentation chains and the accepted fragmentation chain form a fragmentation history [^7] For the accepted fragmentation chain, finalTwo proceeds to generate two additional hadrons from the remaining string. Thus, the total number of hadrons will be the number of string breaks contained in the accepted chain plus the additional two hadrons, which are not counted as string breaks in this work. This is because their kinematic distribution does not depend on the $a$ and $b$ Lund parameters as long as the low-energy threshold that determines whether finalTwo is applied remains fixed, and thus do not need to be reweighted for the example we consider here.,
@@
Explicitly, $e_{h}$ is an unordered list of $N_{\text{had}}=N_{h,N_{h}}+2$ laboratory frame four momenta, $(E_{i},\vec{p}_{i})$, and masses, $m_{i}$, of the produced hadrons [^8] Additional information could be optionally included for each hadron such as flavor composition or charge.,
@@ 2.2 Details about the Homer method
In the Lund string fragmentation model the probability distribution for a given string break $\vec{s}_{hcb}$ depends on the $\rm{}^{\phantom{,}\text{string}}$ of the string fragment that is emitting the hadron. For convenience, we will write this conditional probability as
@@
Homer aims to learn a fragmentation weight $w_{\text{s}}^{\text{data}}(\vec{s}_{hcb})$ for individual hadron emissions. This is achieved by taking an initial guess for the fragmentation probability function $p_{\text{sim}}(\vec{s}_{hcb})$ from the baseline simulation model, e.g.Pythia, and reweight it to a data-driven fragmentation function
@@
In this paper we limit ourselves to emissions of pions for simplicity. The baseline simulation model is taken as eq. 3, with parameters chosen to differ sufficiently from those used to produce the synthetic data. This is used to generate $N_{\text{sim}}$ baseline simulation model events. The baseline simulation model events are then compared with the experimental run dataset that contains $N_{\text{data}}$ events. In obtaining $w_{\text{s}}^{\text{infer}}(\vec{s}_{hcb})$, Homer only uses information that is experimentally accessible. The structure of the hadronization function is simplified due to the assumption of only including pions, which results in a simpler mass and transverse mass spectrum. This reduces the scope of the problem from a training perspective. The exploration of more realistic scenarios, including data-driven flavor selection is left for future work.
@@
and thus is a known analytic function, up to the normalization constants for each $f(z)$. The exact weights for an event, $w_{\text{exact}}(e_{h})$, are then built from the exact single emission weights $w_{\text{s}}^{\text{exact}}(\vec{s}_{hcb})$.
@@ 2.2.1 Step 1: event classifier
where $y(\vec{x}_{h})\in[0,1]$ is the output of a classifier for the input vector $\vec{x}_{h}$, and $N_{\text{sim}}$ and $N_{\text{data}}$ are the number of training samples per class which we assume to be equal in eq. 14. If needed, we reweigh the classes to ensure that each class has the same weighted number of events. This choice of the loss function guarantees that when the classifier training converges, we can obtain a good estimator for the event weight of eq. 11 from the output of the classifier,
@@
Here, the classifier input is still the vectors of observables for each event, $\vec{x}_{h}$ of eq. 13. However, measured data from experiments is oftentimes not available on an event by event basis, and instead only published in the form of histograms for each observable $\mathcal{O}_{i}$, which are assigned to $n_{i}$ bins. The loss function on which the classifier is trained is constructed from the binned measured data and the reweighted baseline simulation model. Concretely, the loss function is given by
@@
For this scenario we represent an event $e_{h}$ as a vector $\vec{x}_{h}$ of four momenta, see eq. 8, which contains all the information available in the event expect particle type and charge. This is the limiting case of the unbinned scenario and therefore proceeds analogously; the available information about the experimental run is represented by a collection of all $\vec{x}_{h}$, giving the input vector $\vec{X}_{\text{data}}=\{\vec{x}_{1},\ldots,\vec{x}_{N_{\text{data}}}\}$. The $N_{\text{sim}}$ baseline simulation model events are collected in the baseline simulation model input vector, $\vec{X}_{\text{sim}}=\{\vec{x}_{1},\ldots,\vec{x}_{N_{\text{sim}}}\}$. As in the unbinned scenario, a classifier is trained using the standard BCE loss function of eq. 14, and we obtain $w_{\text{class}}(e_{h})\approx w_{\text{exact}}(e_{h})$ through eq. 15.
@@ 2.2.2 Step 2: inference of fragmentation weights
The goal of this step is to construct the appropriate single emission weights $w_{\text{s}}$ such that the probability for each string break that produces the emission, $p_{\text{data}}(\vec{s}_{hcb})$ of eq. 10, will reproduce data by reweighting the baseline simulation model string breaks. However, there are two complications. First, in Pythia, the baseline fragmentation history also contains the string fragmentation chains that do not pass the finalTwo filter. That is, the finalTwo filter divides the fragmentation chains $\{\vec{S}_{hc}\}$ in two: chains that pass the finalTwo filter, a set consisting of $\{\vec{S}_{hc}^{\text{acc}}\}$, and those that do not, set $\{\vec{S}_{hc}^{\text{rej}}\}$. Second, the measurable event quantities, i.e., the momenta of the outgoing hadrons of $e_{h}$, given by eq. 8, form an unordered list, since there is no information about the sequence of causally disconnected string breaks. This means that the fragmentation chain $\vec{S}_{hN_{h}}$ and a different fragmentation chain, $\vec{S}_{hN_{h}}^{\prime}$, which gives rise to exactly the same hadron four momenta except with a different order of emissions, are indistinguishable physically. In general this does not correspond to simply permuting the entries in $\vec{S}_{hN_{h}}$, since the required boosts to the laboratory frame depend on previous hadron emissions. That is, the two fragmentation chains give rise to the same observable event, $e_{h}=e(\vec{S}_{hN_{h}})=e(\vec{S}_{hN_{h}}^{\prime})$.
@@
where the summation over $N_{h}$ in eq. 18 counts the number of fragmentation chains that are rejected in the simulation, before $\vec{S}_{hN_{h}}$ is accepted. [^9] Note that the $N_{h}$ index for $\vec{S}_{hN_{h}}$ in eq. 18 is a dummy index, and thus the $p_{\text{rej}}$ is truly independent from the second term in eq. 18. Similarly, the total probability for the accepted fragmentation chains is given by
@@
Above, the probabilities for individual fragmentation chains are products of string breaks,
@@
We label the probabilities for this modified model with the “data” label and the equivalent probabilities of eqs. 18, 19, 20 and 21 as $p_{\text{data}}(e_{h})$, $p_{\text{data}}^{\text{rej}}$, $p_{\text{data}}^{\text{acc}}$ and $p_{\text{data}}(\vec{S}_{hc})$, respectively. With real measured data, these probabilities may only be approximate, but for the synthetic data that we use as an example in this paper, we know these exact probabilities must exist. Below, we describe how the best estimate for $p_{\text{data}}(\vec{s}_{hcb})$ is found.
@@
The neural network of eq. 27 takes as the input the seven-dimensional string break vector, $\vec{s}_{hcb}$ given by eq. 4, and outputs the weight $w_{\text{s}}^{\text{infer}}$ for this string break. Since the event weight, $w_{\text{infer}}(e_{h},\theta)$, involves products of multiple weights, see eq. 25, it is easier to learn the logarithm of $w_{\text{s}}$. In fact, it is numerically expedient to introduce $\ln g_{\theta}(\vec{s})$ as a difference of two neutral networks
@@
where $\mathcal{L}_{C}$, given in eq. 34 below, ensures that the event-level weights $w_{\text{infer}}(e_{h},\theta)$ reproduce the weights $w_{\text{class}}(e_{h})$ of eq. 11 well, which were learned in Step 1. The second term, $\mathcal{L}_{12}$, given in eq. 36 below, is a regularization term that ensures a proper convergence of the $g_{1}$ and $g_{2}$ NNs towards a solution that satisfies the conditional structure imposed by the string breaks within the Lund string model. In the remainder of this section we motivate the forms of these two loss functions.
@@ 2.2.3 Step 3: Homer output
More importantly, as we have shown, the expectation values that enter the calculation of event weights, can be estimated efficiently using simulated Monte Carlo samples where we have access to the simulated histories. This allows us to accurately estimate a new fragmentation function without explicit access to the baseline fragmentation function. We emphasize that we are taking advantage of the fact that for simulated events we have both the underlying fragmentation history, as produced by the model, and observable quantities that can be compared with measurements. If we learn how to reweight a given history in order to match the measured event distribution, we are effectively updating the fragmentation model.
@@ 2.3 Contrasting the Homer method with GANs
Upon initial inspection, the Homer method may appear similar in spirit to the adversarial strategies employed when training GANs[3, 4]. However, the optimization sequence in the Homer method differs significantly from GANs. Rather than incorporating the weights $w(e_{h})$ into a GAN-like loss function of the form

### v2

@@ 2.1 Lund string fragmentation model for the $q\bar{q}$ case
where the term $bA$ corresponds to the imaginary part of the action of a massless string. The area $A$ is the space-time area of the string scaled by $\kappa^{2}$. Generating the transition from a string state to a hadronic state which fulfills eq. 1, is now a question of selecting a specific implementation. In the Lund model, the increasing separation between the quark and the antiquark, makes it energetically favorable to create $q_{j}\bar{q}_{j}$ pairs out of the vacuum. The string therefore breaks into spacelike separated fragments, i.e., hadrons. Due to the spacelike separation, there is no preferred time-ordering, and the hadronization process could therefore be described either by an inside-out cascade, starting from the center and fragmenting outwards, or an outside-in cascade, starting from the string ends. The Lund model makes the latter selection. Starting from one string end, a string break now produces a hadron containing the string end quark (or anti-quark), and the anti-quark (or quark) from the string break. In this way, the $j$’th string break will produce the hadron containing $q_{j-1}\bar{q}_{j}$ or vice versa. We call this sequence of multiple hadron emissions from a single string fragment a fragmentation chain.
@@
Since hadron masses are discrete and the fragmentation function in eq. 3 carries no information about the global state of the string, care must be taken near the end of the iterative process to produce a physical state. When the remaining string system has an invariant mass below a chosen low-energy threshold, it is then converted into a final pair of hadrons. In Pythia, this conversion is performed by the finalTwo algorithm. The finalTwo method effectively works as a filter by checking whether the final hadrons can be produced on-shell while ensuring that the overall fragmentations follow the left-right symmetric Lund fragmentation function. If this is not the possible, the entire generated fragmentation chain is rejected, and the simulation of hadronization starts anew, taking again the original string as the starting point. This process can be repeated several times until the finalTwo step is successful.
@@
String breaks are the production points of hadrons in the hadronization process. They are described by seven dimensional vectors containing: the light-cone momentum fraction of the hadron, $z$; the two-dimensional momentum kick $\Delta\rm=(\Delta p_{x},\Delta p_{y})$ of the emitted hadron; the hadron mass $m_{ij}$; a boolean that encodes whether the string break occurred at the positive or the negative end of the string, stored within Pythia as fromPos; and the transverse momentum of the string before the breaking, =$\rm{}^{\phantom{,}\text{string}}=(p_{x}^{\text{string}},p_{y}^{\text{string}})$ [^6] In general, the energy of the string in its center-of-mass frame, $E_{\text{string}}$, would also be included in this list. We are able to use the shorter list in eq. 4, since the fragmentations in the Lund string model do not depend explicitly on $E_{\text{string}}$. Additional information must also be included in the case where gluons are attached to the string, and the initial state changes between events.
@@
while a vector of rejected fragmentation chains and the accepted fragmentation chain form a fragmentation history, [^7] For the accepted fragmentation chain, finalTwo proceeds to generate two additional hadrons from the remaining string. Thus, the total number of hadrons will be the number of string breaks contained in the accepted chain plus the additional two hadrons, which are not counted as string breaks in this work. This is because their kinematic distribution does not depend on the $a$ and $b$ Lund parameters as long as the low-energy threshold that determines whether finalTwo is applied remains fixed, and thus do not need to be reweighted for the example we consider here.
@@
Explicitly, $e_{h}$ is an unordered list of $N_{\text{had}}=N_{h,N_{h}}+2$ laboratory frame four momenta, $(E_{i},\vec{p}_{i})$, and masses, $m_{i}$, of the produced hadrons, [^8] Additional information could be optionally included for each hadron such as flavor composition or charge.
@@ 2.2 Details about the Homer method
In the Lund string fragmentation model, the probability of a given string break $\vec{s}_{hcb}$ depends on the $\rm{}^{\phantom{,}\text{string}}$ of the string fragment that is emitting the hadron. We write this conditional probability as
@@
Homer aims to learn a fragmentation weight $w_{\text{s}}^{\text{data}}(\vec{s}_{hcb})$ for individual hadron emissions. This is achieved by taking an initial guess for the fragmentation probability function $p_{\text{sim}}(\vec{s}_{hcb})$ from the baseline simulation model, e.g.Pythia, and reweighting it to a data-driven fragmentation function
@@
In this work, we make the simplifying assumption that only pions are produced in fragmentation, leading to a simpler transverse mass spectrum (the exploration of more realistic scenarios, including data-driven flavor selection is left for future work). The baseline simulation model is taken as eq. 3, with parameters chosen to differ sufficiently from those used to produce the synthetic data. $N_{\text{sim}}$ baseline simulation model events are generated. The baseline simulation model events are then compared with the experimental run dataset that contains $N_{\text{data}}$ events. In obtaining $w_{\text{s}}^{\text{infer}}(\vec{s}_{hcb})$, Homer only uses information that is experimentally accessible.
@@
[…]
@@ 2.2.1 Step 1: event classifier
where $y(\vec{x}_{h})\in[0,1]$ is the output of a classifier for the input vector $\vec{x}_{h}$, and $N_{\text{sim}}$ and $N_{\text{data}}$ are the number of training samples per class, which we assume to be equal in eq. 14. If needed, we weight the classes to ensure that each class has the same weighted number of events. This choice of the loss function guarantees that when the classifier training converges, we can obtain a good estimator for the event weight of eq. 11 from the output of the classifier,
@@
Here, the classifier input is still the vectors of observables for each event, $\vec{x}_{h}$ of eq. 13. However, measured data is partitioned into disjoint bins $n_{i}$ of observable $\mathcal{O}_{i}$ values. The classifier is trained using a loss function constructed from the binned measured data and the reweighted baseline simulation model,
@@
For this scenario we represent an event $e_{h}$ as a vector $\vec{x}_{h}$ of four momenta, see eq. 8. $\vec{x}_{h}$ contains all the information available in the event expect particle type and charge. This is the limiting case of the unbinned scenario and therefore proceeds analogously; the available information about the experimental run is represented by a collection of all $\vec{x}_{h}$, giving the input vector $\vec{X}_{\text{data}}=\{\vec{x}_{1},\ldots,\vec{x}_{N_{\text{data}}}\}$. The $N_{\text{sim}}$ baseline simulation model events are collected in the baseline simulation model input vector, $\vec{X}_{\text{sim}}=\{\vec{x}_{1},\ldots,\vec{x}_{N_{\text{sim}}}\}$. As in the unbinned scenario, a classifier is trained using the standard BCE loss function of eq. 14, and we obtain $w_{\text{class}}(e_{h})\approx w_{\text{exact}}(e_{h})$ through eq. 15.
@@ 2.2.2 Step 2: inference of fragmentation weights
The goal of this step is to construct the appropriate single emission weights $w_{\text{s}}$ such that the probability for each string break that produces the emission, $p_{\text{data}}(\vec{s}_{hcb})$ of eq. 10, will reproduce data by reweighting the baseline simulation model string breaks. However, there are two complications. First, in Pythia, the baseline fragmentation history also contains the string fragmentation chains that do not pass the finalTwo filter (as explained in section 2.1). That is, the finalTwo filter divides the fragmentation chains $\{\vec{S}_{hc}\}$ in two: chains that pass the finalTwo filter, a set consisting of $\{\vec{S}_{hc}^{\text{acc}}\}$, and those that do not, set $\{\vec{S}_{hc}^{\text{rej}}\}$. Second, the measurable event quantities, i.e., the momenta of the outgoing hadrons of $e_{h}$, given by eq. 8, form an unordered list, since there is no information about the sequence of causally disconnected string breaks. This means that two fragmentation chains, $\vec{S}_{hN_{h}}$ and $\vec{S}_{hN_{h}}^{\prime}$, which give rise to exactly the same hadron four momenta except with a different order of emissions, are physically indistinguishable. That is, the two fragmentation chains give rise to the same observable event, $e_{h}=e(\vec{S}_{hN_{h}})=e(\vec{S}_{hN_{h}}^{\prime})$.
@@
where the summation over $N_{h}$ in eq. 18 counts the number of fragmentation chains that are rejected in the simulation before $\vec{S}_{hN_{h}}$ is accepted. [^9] Note that the $N_{h}$ index for $\vec{S}_{hN_{h}}$ in eq. 18 is a dummy index, and thus the $p_{\text{rej}}$ is truly independent from the second term in eq. 18. Similarly, the total probability for the accepted fragmentation chains is given by
@@
Note that the probabilities for individual fragmentation chains are products of string breaks,
@@
We use the label “data” for the probabilities that describe the measured distributions and the equivalent expressions of eqs. 18, 19, 20 and 21, i.e.$p_{\text{data}}(e_{h})$, $p_{\text{data}}^{\text{rej}}$, $p_{\text{data}}^{\text{acc}}$ and $p_{\text{data}}(\vec{S}_{hc})$, respectively. With real measured data, these probabilities may only be approximate, but for the synthetic data that we use as an example in this paper, we know these exact probabilities must exist. Below, we describe how the best estimate for $p_{\text{data}}(\vec{s}_{hcb})$ is found.
@@
The neural network of eq. 27 takes as input the seven-dimensional string break vector, $\vec{s}_{hcb}$ given by eq. 4, and outputs the weight $w_{\text{s}}^{\text{infer}}$ for this string break. Since the event weight, $w_{\text{infer}}(e_{h},\theta)$, involves products of multiple weights, see eq. 25, it is easier to learn the logarithm of $w_{\text{s}}$. In fact, it is numerically expedient to introduce $\ln g_{\theta}(\vec{s})$ as a difference of two neutral networks
@@
where $\mathcal{L}_{C}$, given in eq. 34 below, ensures that the event-level weights $w_{\text{infer}}(e_{h},\theta)$ reproduce the weights $w_{\text{class}}(e_{h})$ of eq. 11 well, which were learned in Step 1. The second term, $\mathcal{L}_{12}$, given in eq. 36 below, is a regularization term that ensures a proper convergence of the $g_{1}$ and $g_{2}$ NNs towards a solution that satisfies the conditional structure imposed by the string breaks within the Lund string model. In the remainder of this section, we motivate the forms of these two loss functions.
@@ 2.2.3 Step 3: Homer output
More importantly, as we have shown, the expectation values that enter the calculation of event weights can be estimated efficiently using simulated Monte Carlo samples where we have access to the simulated histories. This allows us to accurately estimate a new fragmentation function without explicit access to the baseline fragmentation function. We emphasize that we are exploiting the fact that, for simulated events, we have both the underlying fragmentation history, as produced by the model, and observable quantities that can be compared with measurements. If we learn how to reweight a given history to match the measured event distribution, we are effectively updating the fragmentation model.
@@ 2.3 Contrasting the Homer method with GANs
Upon initial inspection, the Homer method may appear similar in spirit to the adversarial strategies employed when training GANs[3, 4]. However, the optimization sequence in the Homer method is significantly different. Rather than incorporating the weights $w(e_{h})$ into a GAN-like loss function of the form